OpenStudy (anonymous):

A pool company forgets to bring their logarithmic charts, but they need to raise the amount of hydronium ions in a pool by 0.50. Using complete sentences, explain how your graph can be used to solve 10–y = 0.50. Find the approximate solution. @agent0smith

OpenStudy (anonymous):

@ikram002p @nincompoop

OpenStudy (ikram002p):

r u sure from this ?10–y = 0.50 isnt it 10^–y = 0.50??

OpenStudy (anonymous):

you're right

ganeshie8 (ganeshie8):

... how your graph can be used to solve... where is the graph ? :)

OpenStudy (ikram002p):

so its like this 10^-y=1/10^y  10^y= 2 take the log y=log 2